A group of 7 women and 7 men must select a four-person committee. How many committees are possible if it must consist of the following?
two women and two men?
a majority of women?

Answer by sudhanshu_kmr(1152)


A) when 2 women and 2 men
no. of ways = 7C2 * 7C2 = 21 *21 = 441

B) a majority of women..
possible cases are...
4 women, 0 men
3 women, 1 men
no. of ways = 7C4* 7C0 + 7C3* 7C1 = 35 + 35*7 = 280